# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ 4.2.2. Ledger State Update
Updating the ledger is another operation that may reveal sensitive information to other parties. We review the approaches for account-based and token-based ledgers.
@@
With unspent tokens, the payment operation simultaneously invalidates previous tokens and generate new ones. The state of the system is represented by the list of unspent tokens999Since blockchains store transactions in an append-only log, they trace the derivation history of each unspent token. This helps implement a decentralize yet scalable system.. Unspent tokens cannot be easily marked as spent without compromising sender’s anonymity. Most of the approaches, including CryptoNote (Van Saberhagen, 2013), Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), and their extensions (e.g., (Garman et al., 2017; Androulaki and Karame, 2014; Xue et al., 2023)), let the payer spend a token by using a serial number and a ZKP that indicates that the number is a valid opening to some commitment stored in an unspent token. Monero (Noether et al., 2016) (and its extension Traceable Monero (Li et al., 2021)) uses one-time signatures for unspent tokens, which can be used only once to spend the token. At spending time, the ledger registers a salted hash of the public key. MimbleWimble (and PRCash that reuses its transaction format (Wüst et al., 2019)) eliminates the idea of performing transactions between addresses; instead, it uses a binding and hiding commitment scheme (Pedersen, 1992), together with Bulletproof range proofs (Bünz et al., 2018). By using commitments as outputs, coins act as having their own private key, the knowledge of which, along with the knowledge of their value, would enable spending the funds. This guarantees anonymity and unlinkability.
@@ 4.2.3. Handling Change
Account-based ledgers use transactions to transfer exact value between payer and payee, therefore these kind of systems do not need to support change.
Instead, token-based ledgers propose different approaches for handling change. Systems relying on Chaumian tokens do not easily support change or partial spending (e.g., (Dold, 2019)). Tomescu et al. (Tomescu et al., 2022) allow multiple inputs and outputs for a single transaction, but requires the payer to prove value preservation in ZK. The proof consists in two parts: the first proves that the sum of inputs equals the sum of outputs; the second is a range proof on the output value (needed because ZKPs works with modulo arithmetic and the equality proof is not enough to guarantee value preservation). System based on unspent tokens inherit the change implementation strategy by Bitcoin, where the change is another unspent token owned by a (new) address managed by the payer (e.g., (Miers et al., 2013; Androulaki et al., 2020; Elfadul et al., 2024)) or a commitment controlled by the payer (e.g., (Noether et al., 2016; Wüst et al., 2019)).
@@ 4.2.4. Support for Different Denominations
Chaum (Chaum, 1983) considered a simple system where tokens have implicit value. A few extensions consider the possibility to emit tokens with different denominations (e.g., (Dold, 2019; Tomescu et al., 2022)). For example, Dold (Dold, 2019) in GNU Taler use multiple denomination key pairs for blind-signing coins of different monetary values. Conversely, Tomescu et al. (Tomescu et al., 2022) explicitly represent the token value by concealing it within a commitment. As regards unspent token-based ledgers, Zerocoin does not support arbitrary denominations (Miers et al., 2013). This design simplifies the ZKP used for anonymity but introduces usability limitations, as users must split or combine coins when making payments. Zerocash (Ben Sasson et al., 2014) supports arbitrary denominations by using homomorphic commitments, allowing users to transact any amount while preserving privacy. The remaining approaches allow transacting tokens of different denominations, as each unspent token internally holds its value (or a commitment to it). Among them, CryptoNote (Van Saberhagen, 2013) does not hide transaction amounts, making it vulnerable to statistical analysis—e.g., identifying change as the smaller output. Using multiple change addresses can obscure amounts but leads to issues like proliferation of dust transactions (Noether et al., 2016). Hiding values with commitments prevents such analysis.
@@ 4.2.5. Token Life-cycle
Two main approaches to double spending protection emerge: deposit-time checks and publicly verifiable ledgers. Centralized systems usually relying on Chaumian tokens do not provide a transparent ledger or the primitives for proactive verification of double spending for the payee. Instead, they check token validity at deposit time, i.e., when the recipient interact with the bank to deposit the token (e.g., (Chaum, 1983; Dold, 2019; Tomescu et al., 2022; Goodell et al., 2023)). The public visibility of the ledger in blockchains (e.g., (Nakamoto, 2008)) allows payee to autonomously check whether an received token has been already spent. This validity check is also possible in privacy preserving systems, which usually require the payer to provide a ZKP of token validity (e.g., (Miers et al., 2013; Poelstra, 2016; Noether et al., 2016; Li et al., 2021; Xue et al., 2023; Yu et al., 2024)).
@@
Only a few works (e.g., (Hayes, 1990; Goodell et al., 2023)) explore mechanisms for transferring Chaumian tokens, i.e., enabling safe ownership changes. Hayes (Hayes, 1990) proposes a system where each coin owner signs a chain of transactions to ensure traceability. For example, if user receives a token from the bank, he can pay another user by issuing a token for and attaching the token received from the bank. When the token chain becomes too long, the owner can exchange it for a fresh token via the bank. However, this approach is vulnerable to double spending, which is mitigated by imposing limited token lifetimes. Similarly, Goodell et al. (Goodell et al., 2023) define tokens that carry a proof of provenance: it verifies validity without requiring external interaction. This proof is a list of token commitments including the initial commitment and signature by the issuer.
@@
With Chaumian tokens, the payer-payee interaction takes place off-the-ledger; therefore, the payee should be online to deposit the token, after a validity check by the token issuer. This kind of interaction requires the payee to immediately claim the token to prevent double spending attacks (Chaum, 1983). One-time signatures can be used to reveal misbehaving senders’ identity, enabling revocation and blacklisting (Hayes, 1990). Similarly, in (Camenisch et al., 2006), spending a token multiple times reveals the sender’s identity. Conversely, Sander and Ta-Shma (Sander and Ta-Shma, 1999) avoids relying on immediate claiming of tokens by using Merkle hash trees to publicly authenticate coin validity through inclusion proofs, rather than secret signatures. Coins remain valid as long as users maintain updated hash chains to public tree roots, which are periodically broadcast. This enables offline, asynchronous payments where coins can be spent without real-time interaction with the bank.
@@ 4.3. Privacy Features
We delve into the details of the technical solutions used to obtain the different privacy goals. Privacy in digital currencies is often defined through different sub-properties focusing on anonymity of participants, value confidentiality, plausible deniability, and unlinkability (e.g., (Almashaqbeh and Solomon, 2022; Amarasinghe et al., 2019; Bernal Bernabe et al., 2019; Genkin et al., 2018)).
@@ 4.3.1. Sender Anonymity
Sender anonymity relates to the protection of the payer’s identity, i.e., the transaction sender. Some work implement asymmetric anonymity, where only one party remains anonymous (e.g., (Chaum, 1983; Sander and Ta-Shma, 1999; Miers et al., 2013; Dold, 2019; Zhang et al., 2019; BIS Innovation Hub, 2023; Goodell et al., 2023; Yu et al., 2024; Androulaki and Karame, 2014)). The main techniques for achieving sender anonymity are: blind signatures (e.g., (Chaum, 1983; Dold, 2019; Goodell et al., 2023; Liu et al., 2022)) and their variants—one-time (e.g., (Hayes, 1990)), threshold (e.g., (Sarencheh et al., 2025; Kiayias et al., 2022; Androulaki et al., 2020)), and randomizable (e.g., (Tomescu et al., 2022; Wüst et al., 2019))—, ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), group signatures (e.g., (Zhang et al., 2019)), and commitments with ZKPs (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)).
@@
Chaum (Chaum, 1983) recurs to blind signature to unlink the sender request for signing a token and its later spending. To obtain revocable anonymity and reveal the sender on double spending, Hayes (Hayes, 1990) introduces a one-time blind signature schemes. However, since each transaction requires a new signing key, the storage and computation management is cumbersome; moreover, one-time signatures often require large signature size and expensive verification, making them impractical for high-throughput systems. UTT (Tomescu et al., 2022) realizes blind signatures using randomizable signatures by Pointcheval and Sanders (Pointcheval and Sanders, 2016), which allow signatures to be altered using fresh randomness without compromising their validity. All these schemes rely on a central issuer, who can represent a single point of failure. This can be mitigated, e.g., by using threshold blind signatures, where a quorum of authorities must cooperate to create a valid blind signature (e.g., (Androulaki et al., 2020; Kiayias et al., 2022; Sarencheh et al., 2022)).
@@
Other works reuse the idea of mixing protocols through the concept of ring signatures, which conceals the sender’s public key—hence, his identity—in a group of other users’ public keys (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Li et al., 2021; Elfadul et al., 2024)); this group is referred to as anonymity set. For example, CryptoNote (Van Saberhagen, 2013) uses one-time (linkable) ring signatures for sender anonymity. For each transaction, the sender generates a one-time key and derives a key image, then selects an anonymity set and produces a ring signature including the key image and a verification challenge. The key image is stored on a public ledger managed by a centralized entity to detect double-spending. Verifying a signature requires the public keys of all users in the ring; thus, larger anonymity sets increase privacy but also transaction size. Different privacy-preserving cryptocurrency systems use ring signatures for obtaining sender’s anonymity, including Monero101010Möser et al. (Möser et al., 2018) showed weaknesses in Monero’s key sampling: of transactions do not include any other public key, sampling is not really random, and real inputs can be inferred with accuracy. (Noether et al., 2016), Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024). However, studies show that Monero’s anonymity can be compromised through transaction analysis (e.g., (Möser et al., 2018; Kappos et al., 2018; Kumar et al., 2017)).
@@
Considering the requirement of revoking sender’s anonymity if needed, Zhang et al. (Zhang et al., 2019) use group signatures where a trusted centralized entity sets up the anonymity set and can, optionally, uncover the senders’ identity for inspection. Differently from ring signatures where the anonymity set is dynamic and can change for each transaction, group signatures have a less flexible management.
@@
To be valid, unspent tokens require proving their ownership. To implement sender’s anonymity, existing approaches rely on two key tricks. First, the unspent token includes a commitment to its serial number; the serial number is revealed only to mark the token as spent (see Section 4.2.5). Second, the spending operation includes a ZKP that demonstrates knowledge of the opening to a commitment of an unspent coin. Sander and Ta-Shma (Sander and Ta-Shma, 1999) use multiple Merkle trees to track valid Chaumian tokens. To prove validity, the sender provides a NIZK proof showing that he knows a valid coin in the tree and that the coin’s serial number and a hash chain correctly lead to one of the live Merkle tree roots. Zerocoin (Miers et al., 2013), and its extensions (e.g., (Ben Sasson et al., 2014; Garman et al., 2017)), adopts ZKPs as well. In these systems, the payer publishes a transaction that includes the recipient’s address, an empty origin address, a ZKP, and the commitment opening. The ZKP prevents linking the sender to any specific commitment; the commitment itself hides the opening using a random value that is never revealed, preventing any association between the two.
@@
Privacy-preserving account-based models mostly rely on ZKPs to obtain sender’s anonymity (e.g., (Gross et al., 2021; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024)). Specifically, a commitment to the account state is used to trace the balance (along with other information, as detailed later); account state updates include ZKPs to prove correctness. For example, Gross et al. (Gross et al., 2021) use the ideas of Zerocash by committing on the accounts state and by invalidating them on update. The authors suggest using Merkle trees to prove that a transaction proposal refers to an existing commitment in the ledger without pointing to (and thus revealing) it. Platypus (Wüst et al., 2022) uses a model where payer and payee exchange with the central bank commitments to their updated accounts, the transaction amount, a serial number, and two ZKPs proving the update’s validity and its link to a prior certified state. The payer sends his part to the payee, who completes it and interacts with the central bank. If the serial and the ZKPs are valid, the bank updates a publicly available log. Conversely, PEReDi (Kiayias et al., 2022) achieves sender anonymity through anonymous channels, an encrypted ledger, and a threshold signature scheme. Specifically, the sender and receiver each submit partial (specular) transactions that update only their respective account states. These transactions are encrypted using threshold encryption. Sender and receiver must both actively engage in updating their accounts. Solidus (Cecchetti et al., 2017) uses banks as proxies; therefore, transactions are not explicitly exchanged between users but among their respective banks.
@@ 4.3.2. Receiver Anonymity
The key techniques for obtaining receiver’s anonymity are one-time destination key or stealth address (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), encrypted address (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)), commitments as output (e.g., (Poelstra, 2016; Wüst et al., 2019)), and updatable public key (e.g., (Fauzi et al., 2019; Papadoulis et al., 2024)). Obtaining receiver anonymity with Chaumian tokens is challenging, as the payee must interact with the bank to deposit the received tokens. Hayes (Hayes, 1990) suggests using one-time aliases to obtain privacy, even though the idea is not detailed for practical usages. A similar idea can be found in CryptoNote (Van Saberhagen, 2013), where the sender generates a one-time destination key by combining a random number with the receiver’s public key—essentially performing a Diffie-Hellman exchange to derive a shared secret. Because this process requires two elliptic curve keys, a CryptoNote address is twice the size of a standard elliptic curve public key. The transaction is then sent to the one-time public key and includes a commitment to the random value. The receiver scans all the incoming transaction using his private key, reconstructs the one-time destination public key from the commitment, and checks if it matches; if it does, he recognizes the payment as intended for him. The sender and receiver never interact directly. The most obvious disadvantage of this approach is that every receiver has to read every transaction to identify those belonging to him. This approach has been later integrated in other privacy-preserving cryptocurrencies, including Monero (Noether et al., 2016), which calls the one-time receiver address as stealth address, Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024) (see Section 3.6.1).
@@
In Zerocash (Ben Sasson et al., 2014) (unspent token-based ledger), the sender includes an encrypted version of the recipient’s address in the transaction (using asymmetric encryption). Therefore, periodically, each user has to scan the ledger, decrypt transactions using his private key, and identify those intended for him. Similarly, UTT (Tomescu et al., 2022) (Chaumian token-based system) ensures receiver’s anonymity using homomorphic commitments, randomizable signatures, and anonymous (Boneh-Franklin) IBE scheme (Boneh and Franklin, 2003). A coin is a commitment to a tuple containing the owner’s identifier, a serial number issued by the central bank, and its value. To be valid, each coin comes with a (randomizable) signature by the central bank. To make a payment, the sender creates a transaction for the payee, which includes the receiver’s identity commitment, the output value commitment, the encrypted coin details (using the recipient’s key), and ZKPs for correctness. The transaction is sent to the bank for validation. If valid, the bank blindly issues new coins for the recipient and records them on a public ledger—without learning identities or amounts, as they remain hidden via homomorphic commitments111111The homomorphic properties of the commitment scheme allows to directly combine the value commitment, identity commitment, and serial number commitment in a single coin commitment.. The recipient periodically scans the ledger for transactions encrypted for him, decrypts payments with his private key, verifies the amounts, and retrieves the coins’ details. Then, he can randomize the coins before adding them to his wallet.
@@
The unspent token-based ledger MimbleWimble (Poelstra, 2016) (and PRCash (Wüst et al., 2019)) performs transactions between commitments rather than addresses; this guarantees receiver’s anonymity and unlinkability.
@@
Account-based approaches typically achieve receiver anonymity using the same technique as for sender anonymity (e.g., (Wüst et al., 2022; Kiayias et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Lin et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Sarencheh et al., 2025; Papadoulis et al., 2024; Fauzi et al., 2019)). This is done by updating account state commitments and employing ZKPs to verify correctness without revealing the receiver’s identity. We mention Quisquis (Fauzi et al., 2019), where an account is a pair of public key and a homomorphic commitment to its balance. A transaction updates both the public key and the commitment using updatable public key (see Section 3.6.2). Senders select public keys—including the recipient’s and decoys—to form the transaction input, updating only the sender’s and recipient’s balances while leaving others unchanged. Since each address appears at most twice on the ledger, de-anonymization is harder. However, this approach requires users to periodically scan the ledger to update their keys, thus introducing some transaction serialization.

### v2

@@ 4.2.2. Ledger State Update
Updating the ledger may reveal sensitive information to other parties.
@@
With unspent tokens, the payment operation simultaneously invalidates previous tokens and generate new ones. The state of the system is represented by the list of unspent tokens999Since blockchains store transactions in an append-only log, they trace the derivation history of each unspent token. This helps implement a decentralize yet scalable system.. Unspent tokens cannot be easily marked as spent without compromising sender anonymity. Most of the approaches, including CryptoNote (Van Saberhagen, 2013), Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), and their extensions (e.g., (Garman et al., 2017; Androulaki and Karame, 2014; Xue et al., 2023)), let the payer spend a token by using a serial number and a ZKP that indicates that the number is a valid opening to some commitment stored in an unspent token. Monero (Noether et al., 2016) (and its extension Traceable Monero (Li et al., 2021)) uses one-time signatures for unspent tokens, which can be used only once to spend the token. At spending time, the ledger registers a salted hash of the public key. MimbleWimble (and PRCash that reuses its transaction format (Wüst et al., 2019)) uses a binding and hiding commitment scheme (Pedersen, 1992), together with Bulletproof range proofs (Bünz et al., 2018), to define the recipient. By using commitments as outputs, coins act as having their own private key, the knowledge of which, along with the knowledge of their value, enable spending the token. This guarantees anonymity and unlinkability.
@@ 4.2.3. Handling Change
Account-based ledgers use transactions to transfer exact value between payer and payee, therefore these kind of systems do not need to support change. Instead, token-based ledgers propose different approaches for handling change. Chaumian tokens do not easily support change or partial spending (e.g., (Dold, 2019)). Tomescu et al. (Tomescu et al., 2022) allow multiple inputs and outputs for a single transaction, but requires the payer to prove value preservation in ZK. The proof consists in two parts: the first proves that the sum of inputs equals the sum of outputs; the second is a range proof on the output value (needed because ZKPs works with modulo arithmetic and the equality proof is not enough to guarantee value preservation). System based on unspent tokens inherit the change implementation strategy by Bitcoin, where the change is another unspent token owned by a (new) address managed by the payer (e.g., (Miers et al., 2013; Androulaki et al., 2020; Elfadul et al., 2024)) or a commitment controlled by the payer (e.g., (Noether et al., 2016; Wüst et al., 2019)).
@@ 4.2.4. Support for Different Denominations
Chaum (Chaum, 1983) considered a simple system where tokens have implicit value. A few extensions consider the possibility to emit tokens with different denominations (e.g., (Dold, 2019; Tomescu et al., 2022)). For example, Dold (Dold, 2019) in GNU Taler use multiple denomination key pairs for blind-signing coins of different monetary values. Conversely, Tomescu et al. (Tomescu et al., 2022) explicitly represent the token value by concealing it within a commitment. As regards unspent token-based ledgers, Zerocoin does not support arbitrary denominations (Miers et al., 2013). This design simplifies the ZKP used for anonymity but introduces usability limitations, as users must split or combine coins when making payments. Zerocash (Ben Sasson et al., 2014) supports arbitrary denominations by using homomorphic commitments, allowing users to transact any amount while preserving privacy. The remaining approaches allow tokens of different denominations, as each unspent token internally holds its value (or a commitment to it). Among them, CryptoNote (Van Saberhagen, 2013) does not hide transaction amounts, making it vulnerable to statistical analysis—e.g., identifying change as the smaller output. Using multiple change addresses can obscure amounts but leads to proliferation of dust transactions (Noether et al., 2016). Hiding values with commitments prevents such analysis.
@@ 4.2.5. Token Life-cycle
Two approaches protect from double spending: deposit-time checks and publicly verifiable ledgers. Centralized systems with Chaumian tokens do not usually rely on a transparent ledger; instead, they check token validity at deposit time (e.g., (Chaum, 1983; Dold, 2019; Tomescu et al., 2022; Goodell et al., 2023)). The public visibility of the ledger in blockchains (e.g., (Nakamoto, 2008)) allows payee to autonomously check token validity. This is also possible in privacy preserving systems, which usually require the payer to provide a ZKP of token validity (e.g., (Miers et al., 2013; Poelstra, 2016; Noether et al., 2016; Li et al., 2021; Xue et al., 2023; Yu et al., 2024)).
@@
Few works (e.g.,(Hayes, 1990; Goodell et al., 2023)) propose mechanisms for transferring Chaumian tokens. Hayes (Hayes, 1990) proposed chaining transactions, where each owner signs over the previous link; long chains can be refreshed by the bank, though the scheme risks double spending, mitigated by token lifetimes. Goodell et al. (Goodell et al., 2023) instead attach a proof of provenance—a list of commitments with the issuer’s signature—allowing validity checks without external interaction.
@@
With Chaumian tokens, transactions occur off-ledger, so payees must deposit tokens promptly to prevent double spending (Chaum, 1983). Misbehavior can be deterred by one-time signatures that reveal the sender’s identity (Hayes, 1990; Camenisch et al., 2006). Sander and Ta-Shma (Sander and Ta-Shma, 1999) instead use Merkle trees: coin validity is proven via inclusion proofs to public roots, enabling offline asynchronous payments without real-time bank interaction.
@@ 4.3. Privacy Features
Privacy in digital payment systems is often defined through different sub-properties focusing on anonymity of participants, value confidentiality, plausible deniability, and unlinkability (e.g., (Almashaqbeh and Solomon, 2022; Amarasinghe et al., 2019; Bernal Bernabe et al., 2019; Genkin et al., 2018)).
@@ 4.3.1. Sender Anonymity
Asymmetric anonymity indicates that either the sender’s or receiver’s identity is protected (e.g., (Chaum, 1983; Sander and Ta-Shma, 1999; Miers et al., 2013; Dold, 2019; Zhang et al., 2019; BIS Innovation Hub, 2023; Goodell et al., 2023; Yu et al., 2024; Androulaki and Karame, 2014)). Sender anonymity focuses only on hiding the payer’s identity, usually using the following techniques: blind signatures (e.g., (Chaum, 1983; Dold, 2019; Goodell et al., 2023; Liu et al., 2022)) and their variants—one-time (e.g., (Hayes, 1990)), threshold (e.g., (Sarencheh et al., 2025; Kiayias et al., 2022; Androulaki et al., 2020)), and randomizable (e.g., (Tomescu et al., 2022; Wüst et al., 2019))—, ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), group signatures (e.g., (Zhang et al., 2019)), and commitments with ZKPs (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)).
@@
Chaum (Chaum, 1983) proposes blind signature to unlink the sender request of token issuance and its later spending. To reveal the sender on double spending, Hayes (Hayes, 1990) introduces one-time blind signatures. However, since each transaction requires a new signing key, the storage and computation management is cumbersome; moreover, one-time signatures often require large signature size and expensive verification, making them impractical for high-throughput systems. UTT (Tomescu et al., 2022) implements blind signatures using randomizable signatures (Pointcheval and Sanders, 2016), which allow signatures to be altered using fresh randomness without compromising their validity. All these schemes rely on a central issuer, who can represent a single point of failure. This can be mitigated, e.g., by using threshold blind signatures, where a quorum of authorities must cooperate to create a valid signature (e.g., (Androulaki et al., 2020; Kiayias et al., 2022; Sarencheh et al., 2022)).
@@
Other works reuse the idea of mixing through ring signatures, which conceals the sender’s public key in a group of other public keys—the anonymity set (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Li et al., 2021; Elfadul et al., 2024)). For example, CryptoNote (Van Saberhagen, 2013) employs one-time ring signatures: each transaction uses a fresh key and produces a key image, enabling detection of double spends. The signature hides the actual signer among a group, with anonymity scaling with group size; however, larger rings increase transaction size. Different privacy-preserving crypto-asset systems use ring signatures for obtaining sender anonymity, including Monero101010Möser et al. (Möser et al., 2018) showed weaknesses in Monero’s key sampling: of transactions do not include any other public key, sampling is not really random, and real inputs can be inferred with accuracy. (Noether et al., 2016), Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024). However, studies show that Monero’s anonymity can be compromised through transaction analysis (e.g., (Möser et al., 2018; Kappos et al., 2018; Kumar et al., 2017)).
@@
To revoke sender anonymity if needed, Zhang et al. (Zhang et al., 2019) use group signatures where a trusted centralized entity sets up the anonymity set and can, optionally, uncover the senders’ identity for inspection.
@@
To be valid, unspent tokens require proving their ownership. To implement sender anonymity, existing approaches rely on two key tricks. First, the unspent token includes a commitment to its serial number; the serial number is revealed only to mark the token as spent (see Section 4.2.5). Second, the spending operation includes a ZKP that demonstrates knowledge of the opening to a commitment of an unspent coin. Sander and Ta-Shma (Sander and Ta-Shma, 1999) use multiple Merkle trees to track valid Chaumian tokens. To prove validity, the sender provides a ZKP attesting knowledge of a valid coin in the tree and of the coin’s serial number and a hash chain leading to one of Merkle tree roots. Zerocoin (Miers et al., 2013), and its extensions (e.g., (Ben Sasson et al., 2014; Garman et al., 2017)), adopts ZKPs as well: the payer publishes a transaction that includes the recipient’s address, an empty origin address, a ZKP, and the commitment opening. The ZKP prevents linking the sender to any specific commitment; the commitment itself hides the opening using a random value that is never revealed, preventing any association between the two.
@@
Privacy-preserving account-based models mostly rely on ZKPs to obtain sender anonymity (e.g., (Gross et al., 2021; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024)). Specifically, a commitment to the account state is used, with updates including ZKPs to prove correctness. For example, Gross et al. (Gross et al., 2021) use the ideas of Zerocash by committing on the accounts state and by invalidating them on update. The authors suggest using Merkle trees to prove that a transaction proposal refers to an existing commitment in the ledger without pointing to (and thus revealing) it.111111For a broader discussion of set accumulators—including Merkle trees as a key example—and their role in enhancing anonymity, see (Loporchio et al., 2023). Platypus (Wüst et al., 2022) uses a model where payer and payee collaborate to exchange with the central bank commitments to their updated accounts, the transaction (amount and serial number), and two ZKPs proving the update’s validity and its link to a prior certified state. If valid the bank updates a public ledger. Conversely, PEReDi (Kiayias et al., 2022) achieves sender anonymity through anonymous channels, an encrypted ledger, and a threshold signature scheme. Sender and receiver submit partial (specular) transactions that update only their respective account states, encrypted using threshold encryption. Solidus (Cecchetti et al., 2017) uses banks as proxies: transactions are not explicitly exchanged between users but among their respective banks.
@@ 4.3.2. Receiver Anonymity
The key techniques for receiver anonymity are one-time destination key or stealth address (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), encrypted address (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)), commitments as output (e.g., (Poelstra, 2016; Wüst et al., 2019)), and updatable public key (e.g., (Fauzi et al., 2019; Papadoulis et al., 2024)). Obtaining receiver anonymity with Chaumian tokens is challenging, as the payee must interact with the bank to deposit the received tokens. Hayes (Hayes, 1990) suggests using one-time aliases to obtain privacy, even though the idea is not detailed for practical usages. A similar idea can be found in CryptoNote (Van Saberhagen, 2013), where the sender generates a one-time destination key by combining a random number with the receiver’s public key—essentially performing a Diffie-Hellman exchange to derive a shared secret.121212Because this process requires two elliptic curve keys, a CryptoNote address is twice the size of a standard elliptic curve public key. The sender posts the transaction to a one-time public key with a commitment to a random value. The receiver scans all transactions with his private key, reconstructs the destination key, and accepts the payment if it matches. Sender and receiver never interact directly, but each receiver must scan the entire ledger to find his payments. This approach has been later integrated in other systems, including Monero (Noether et al., 2016), which calls the one-time receiver address as stealth address, Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024) (see Section 3.6).
@@
In Zerocash (Ben Sasson et al., 2014) (unspent token-based ledger), the sender includes an encrypted version of the recipient’s address in the transaction (using asymmetric encryption). Therefore, periodically, each user has to scan the ledger, decrypt transactions using his private key, and identify those intended for him. UTT (Tomescu et al., 2022) (Chaumian token-based system) ensures receiver anonymity using homomorphic commitments and anonymous IBE scheme (Boneh and Franklin, 2003). A coin is a commitment to a tuple containing the owner’s identifier, a serial number issued by the central bank, and its value. To be valid, each coin comes with a (randomizable) signature by the central bank. To make a payment, the sender creates a transaction for the payee, which includes the receiver’s identity commitment, the output value commitment, the encrypted coin details (with the recipient’s key), and ZKPs. The transaction is sent to the bank for validation; if valid, the bank blindly issues new coins for the recipient and records them on a public ledger131313The homomorphic properties of the commitment scheme ensures that the bank does not learn identities or amounts, and allows to directly combine the value commitment, identity commitment, and serial number commitment in a single coin commitment.. Recipients periodically scan the ledger to get their coins, to be randomized before adding to the wallet.
@@
The unspent token-based ledger MimbleWimble (Poelstra, 2016) (and PRCash (Wüst et al., 2019)) performs transactions between commitments rather than addresses; this guarantees receiver anonymity and unlinkability.
@@
Account-based approaches typically achieve receiver anonymity using the same technique as for sender anonymity (e.g., (Wüst et al., 2022; Kiayias et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Lin et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Sarencheh et al., 2025; Papadoulis et al., 2024; Fauzi et al., 2019)). This is done by updating account state commitments and employing ZKPs to verify correctness. We mention Quisquis (Fauzi et al., 2019), where an account is a pair of public key and a homomorphic commitment to its balance. A transaction updates both the public key and the commitment using updatable public key (see Section 3.6). Senders select public keys—including the recipient’s and decoys—to form the transaction input, updating only the sender’s and recipient’s balances while leaving others unchanged. Since each address appears at most twice on the ledger, de-anonymization is harder. However, this approach requires users to periodically scan the ledger to update their keys, thus introducing some transaction serialization.